a particle is fired with a speed of 2kmph. the speed with which it will move in interstellar space (escape velocity 8xroot20

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,

By the conservation of energy,
mv2/2 = m((2/60x60)2-ve2)/2    (Where ve= escape velocity = 8√20 km/s)
=>v=√((2/3600)^2-(8√20)^2 )
=>v= which is coming complex
Hence the values given are absurd.
Specify the values and units correctly.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)